### v1

# Self-supervised Learning of LiDAR 3D Point Clouds via 2D-3D Neural Calibration

Abstract
This paper introduces a novel self-supervised learning framework for enhancing 3D perception in autonomous driving scenes. Specifically, our approach, named NCLR, focuses on 2D-3D neural calibration, a novel pretext task that estimates the rigid transformation aligning camera and LiDAR coordinate systems. First, we propose the learnable transformation alignment to bridge the domain gap between image and point cloud data, converting features into a unified representation space for effective comparison and matching. Second, we identify the overlapping area between the image and point cloud with the fused features. Third, we establish dense 2D-3D correspondences to estimate the rigid transformation. The framework not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, understanding their relative pose. We demonstrate NCLR’s efficacy by applying the pre-trained backbone to downstream tasks, such as LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation. Comprehensive experiments on various datasets illustrate the superiority of NCLR over existing self-supervised methods. The results confirm that joint learning from different modalities significantly enhances the network’s understanding abilities and effectiveness of learned representation. Code will be available at https://github.com/Eaphan/NCLR.
Index Terms:
Self-supervised Learning, 3D Perception, Cross-modal, Autonomous Driving.
1 Introduction
The LiDAR technology serves as a vital enhancement to 2D cameras by precisely capturing the surroundings of a vehicle, offering robust performance in challenging conditions, including low light, intense sunlight, or glare from approaching headlights. This 3D perception, derived from LiDAR point clouds, is essential for the effective navigation of autonomous vehicles in three-dimensional spaces. While current leading methods are based on extensive labeled datasets, the process of labeling 3D data is exceedingly expensive and time-intensive, given the limited availability of annotation resources. Consequently, there is a growing imperative to utilize unlabeled data. This approach aims to minimize the need for extensive annotation while enhancing the effectiveness of networks trained on a limited amount of labeled data.
An effective strategy to tackle this challenge involves initially pre-training a neural network solely on unannotated data, for example, by employing a pretext task that eliminates the need for manual labeling. Subsequently, this self-supervised pre-trained network can be fine-tuned for specific downstream tasks. Through thorough pre-training, the network acquires initial weights that serve as a beneficial foundation for additional supervised training. As a result, training the network for a particular downstream task generally demands fewer annotations to achieve comparable performance levels to those of a network trained from scratch. The self-supervised approaches have been very successful in 2D (images), even reaching the level of supervised pre-training [1, 2].
A number of self-supervised approaches have also been successful in the field of point cloud data processing. Several methods explore different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduce contrastive losses to capture the geometric and structural nuances of 3D data [8, 9]. Another class of methods takes the temporal correlation as a form of supervision [10, 11, 12]. For example, STRL [11] processes two adjacent frames from 3D point cloud sequences. It transforms the input with spatial data augmentation and learns an invariant representation. And TARL [10] exploit vehicle movement to match objects in consecutive scans. It then trains a model to maximize the segment-level feature similarities of the associated object in different scans, enabling the learning of a consistent representation across time. Reconstruction-based methods [13, 14] apply masked auto-encoding to point cloud and reconstructed points coordinates using the Chamfer distance. ALSO [15] proposes to use unsupervised surface reconstruction as a pretext task to train 3D backbones.
Figure 1: Illustration of the proposed 2D-3D neural calibration developed as the pretext task for self-supervised learning of 3D LiDAR point clouds. Our method not only learns the local matching from points to pixels via contrastive losses, but also estimates the (unknown) holistic rigid transformation between camera and LiDAR systems. Besides, we propose learnable transformation alignment to fill the domain gap between the image and the point cloud.
Although multi-modal self-supervised learning holds significant promise for point cloud applications, its potential has not been completely realized. Current approaches predominantly utilize unlabeled data from synchronized and calibrated camera-LiDAR setups to pretrain 3D backbones. The core strategy involves identifying corresponding point-pixel pairs and ensuring their representations closely align [16, 17, 18]. However, these methods face two primary limitations: First, existing image-to-point self-supervised learning methods only align the pixels and points locally through contrastive learning, but ignore the holistic spatial relationship between the image and point cloud. Second, these methods neglect the inherent differences in the characteristics of these distinct modalities when aligning the corresponding features of pixels and points. As a result, their performance is still limited.
In this paper, we introduce a novel pretext task, the 2D-3D neural calibration, for the self-supervised pre-training of networks for 3D perception in autonomous driving scenes, as shown in Fig. 1. During pre-training, our pretext task not only learns fine-grained matching from individual points to pixels but also achieves a comprehensive alignment between the image and point cloud data, i.e., understanding their relative pose. Specifically, to impose supervision on the rigid transformation estimation, we propose an end-to-end differentiable framework distinguished by its integration of a soft-matching strategy and a differentiable PnP solver. We posit that this joint learning from different modalities for both local and global-level alignment will enhance the network’s capacity for sophisticated understanding and enable it to develop effective representations. In addition, to fill the substantial domain gap between point clouds and images, we propose a learnable transformation alignment during pre-training, replacing the direct cosine distance alignment. This method converts features into a unified representation space, allowing for more accurate feature comparison and matching beyond the capabilities of cosine distance.
Figure 2: Semantic and panoptic segmentation performance comparison of our method against a recent self-supervised learning method named SegContrast [5] and scratch training on the SemanticKITTI dataset. Our approach achieves comparable results with only 10% annotated data, demonstrating near-equivalent effectiveness to full dataset training from scratch.
To evaluate our method, we carry out comprehensive experiments and compare it against state-of-the-art studies on downstream tasks, including 3D semantic segmentation, object detection, and panoptic segmentation. The results demonstrate that our method outperforms existing self-supervised learning approaches, as evidenced by its superior performance in adapting to various downstream tasks and datasets [19, 20, 21, 22]. Fig. 2 shows that our pre-training method enables the network to reach performance comparable to scratch training with fewer annotations for downstream tasks.
To summarize, the main contributions of this work are as follows:
  • We introduce an innovative perspective to self-supervised learning, centered on achieving a thorough alignment between two distinct modalities.
  • We identify the inherent domain gap between the image and the point cloud and propose the learnable transformation alignment for feature comparison.
  • We propose an end-to-end network featured with the soft-matching strategy and a differentiable PnP solver, achieving state-of-the-art performance on the 2D-3D neural calibration task.
  • We demonstrate the superiority of our method over existing self-supervised learning methods across three distinct downstream tasks in 3D perception.
The rest of this paper is organized as follows. Section 2 provides an overview of related literature pertinent to our research. In Section 3, we present the overall architecture of NCLR and elaborate its principal elements. Section 4 presents an empirical evaluation of our proposed approach across three distinct downstream tasks, along with ablation studies to assess the impact of key components. The paper concludes with Section 5, summarizing our findings.
2 Related Work
2.1 Self-supervised Learning for Images
Self-supervised learning involves training models on unlabeled data by generating pseudo-labels or tasks based on the data itself, without human annotations. The model learns representations by solving pretext problems derived from the inherent structure of the unlabeled inputs. The field of self-supervised learning on images has seen significant evolution over the years. Initial approaches were centered around pretext tasks, which are auxiliary tasks designed to derive meaningful data representations. These tasks involved teaching models to restore color in black-and-white images [23] and to ascertain the spatial relation between image segments [24].
The advent of deep learning brought about a paradigm shift in the field. One such approach is contrastive learning [25, 26]. Contrastive learning involves training networks to distinguish between different augmentations of the same image versus those from distinct images. It minimizes the representation distance for augmentations of the same image (positive pairs) while maximizing it for different images (negative pairs). Generative models [2, 27] have also emerged as a significant trend in self-supervised learning. These models, often employing Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), learned to generate images that closely resemble the training data. By learning to generate images, these models captured the underlying data distribution, which can be leveraged for various downstream tasks. In addition to these, there have been several exciting advancements in the field. Masked image modeling, for instance, involves masking parts of an image and training a model to predict the masked parts [28]. Multi-modal models aim to learn representations across different modalities (e.g., text, image, audio) using self-supervised learning. The Vision Transformer (ViT) adapted the transformer architecture, initially developed for natural language processing, to computer vision tasks [29]. Additionally, there has been explored into incorporating physics-based priors in self-supervised learning.
2.2 Self-supervised Learning for Point Clouds
The advances in self-supervision on point clouds have closely followed the improvements made in images. Early self-supervised methods used pretext tasks such as predicting transformations applied to the point cloud or reconstructing parts of the point cloud [30, 31]. These methods were applied on dense scans of single objects [32, 33].
Discriminative-based Methods. This type of method explored different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduced contrastive losses to capture the geometric and structural nuances of 3D data [8, 9]. Nunes et al. introduced a novel contrastive learning approach focused on understanding scene structure. This method involves extracting class-agnostic segments from point clouds and employing contrastive loss on these segments to differentiate between structurally similar and dissimilar elements [5].
Temporal-consistency-based Methods. STRL [11] utilized pairs of temporally-linked frames from 3D point cloud sequences, applying spatial data augmentation to learn invariant representations. STSSL [12] incorporated spatial-temporal positive pairs, introducing a point-to-cluster technique for spatial object distinction and a cluster-to-cluster method using unsupervised tracking for temporal correlations. TARL [10] used vehicle motion to align objects over time in different scans, focusing on enhancing segment-level feature similarities for consistent temporal representations. These works demonstrated how leveraging temporal information can lead to robust and invariant representations, which are beneficial for various downstream tasks.
Reconstruction-based Methods. Reconstruction-based methods have also been successful for self-supervision on point clouds. Some methods reconstructed point coordinates using the Chamfer distance [13, 14]. Point-BERT randomly masked some patches and fed them into Transformers to recover the original tokens at the masked locations [34]. Recently, ALSO [15] proposed to use unsupervised surface reconstruction as a pretext task to train 3D backbones on automotive LiDAR point clouds. Using the knowledge of occupancy before and after an observed 3D point along a LiDAR ray, it learned to construct an implicit occupancy function and good point features.
Figure 3: The overall pipeline for 2D-3D neural calibration serves as a pretext task for generating self-supervised representation for LiDAR 3D point cloud data. This process involves four key steps: (1) Feature extraction from both the image and point cloud. (2) Detection of overlapping areas using fused features from both modalities. (3) Transformation of image and point cloud features into a unified representation space for similarity computation. We filter the points and pixels in estimated non-overlapping areas. (4) Application of a soft-matching strategy to establish 2D-3D correspondences, followed by the use of a differentiable EPnP solver for camera pose estimation.
Multi-modal Self-supervised Learning Another line of work leveraged synchronized and calibrated cameras and LiDAR to pretrain a 3D backbone. The underlying idea is to find pairs of corresponding points and pixels and ensure that the associated point and pixel representations are as close as possible [16, 17, 18].
2.3 LiDAR-based 3D Perception
LiDAR-based 3D perception is a critical component in many applications, particularly in autonomous driving. It encompasses several tasks, including 3D semantic segmentation, 3D object detection, and 3D panoptic segmentation. Due to space limitations, this paper only discusses the methods we employed. For additional methodologies, we refer readers to the review paper [35].
3D Semantic Segmentation. One of the foundational works in this area is PointNet [36], which marked a significant departure from previous methods by directly processing point clouds for tasks like segmentation. MinkUNet adapted the U-Net architecture for sparse 3D point clouds, using Minkowski Engine for efficient segmentation [37]. SPVCNN combined sparse and continuous convolutions for scalable and efficient 3D point cloud segmentation [38].
3D Object Detection. Advancements in 3D object detection have led to its categorization into voxel-based, point-based, and hybrid methods [39, 40, 41, 42]. Voxel-based techniques convert point clouds into uniform grid structures for feature extraction using sparse convolutions [43, 44]. Point-based methods directly handle raw 3D point clouds to make predictions. Hybrid models, such as PV-RCNN [45] and PDV [46], merge the advantages of both voxel and point-based approaches, optimizing both accuracy and efficiency in detection.
3D Panoptic Segmentation. The task requires a holistic understanding of both objects and scenes. Hong et al. introduced the Dynamic Shifting Network (DS-Net) as a potent framework for panoptic segmentation in the field of point clouds [47]. This approach utilized cylinder convolution tailored for LiDAR data. The features derived were concurrently used by both the semantic and instance branches, functioning through a bottom-up clustering approach.
3 Proposed Method
3.1 Overview
We propose a novel approach to generating self-supervised features for 3D point clouds, utilizing the 2D-3D neural calibration as a pretext task. The 2D-3D neural calibration aims to identify the unknown rigid transformation, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, through a neural network. This transformation aligns the coordinate systems of the camera and the point cloud. We denote the input image as $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ and point cloud as $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}|i=1,2,...,N\}$, where $H$ and $W$ are the height and width of the image, and $N$ is the number of points. To achieve effective representations, our pre-training method not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, i.e., understanding their relative pose. Then the point cloud backbone could be used for downstream 3D perception tasks.
Fig. 3 illustrates our overall framework. Initially, we extract features from both the image and point cloud using dedicated backbones. Next, we identify overlapping regions in both the image and point cloud leveraging the fused features, as overlaps between them are partial. Subsequently, a learnable transformation is applied to harmonize the image and point cloud features into a single representation space before similarity computations between them. The final step involves employing a soft-matching approach to establish 2D-3D correspondences. These correspondences are inputted into a differentiable EPnP solver for camera pose estimation. The proposed framework is designed to be end-to-end trainable.
3.2 Feature Extraction
Both the image and point cloud branches employ an encoder-decoder structure. After the encoder stage, we fuse the downsampled 2D feature map and sampled keypoint features to combine these diverse features into a unified representation.
Transformer-based Feature Fusion. We perform bidirectional feature fusion based on the attention mechanism. We denote the downsampled 2D feature map after encoder as $\mathbf{F}_{E}^{\mathbf{I}}\in\mathbb{R}^{H_{E}\times W_{E}\times C_{E}}$ and features of sampled keypoints as $\mathbf{F}_{E}^{\mathbf{P}}\in\mathbb{R}^{N_{E}\times C_{E}}$. Each fusion layer comprises three components: firstly, a multi-head self-attention layer for image and point features; secondly, a multi-head cross-attention layer that refines each domain’s features using data from the other; and thirdly, a point-wise feed-forward network. These cross-attention layers facilitate the model’s understanding of interrelations and complementary aspects across diverse data types. To incorporate the positional information, we add sinusoidal position encoding to the inputs of transformer layers [48]. Then the fused features are passed to the decoders to obtain higher-resolution feature maps. The outputs of decoder are features $\mathbf{F}^{\mathbf{I}}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$ and $\mathbf{F}^{\mathbf{P}}\in\mathbb{R}^{N\times C}$.
Note that our framework is effective for not only point-based but also voxel-based backbone for the point cloud branch. If we need to pre-train a voxel-based backbone for downstream tasks, we can incorporate it and obtain the final point features by aggregating the voxel-wise features and fine-grained point-wise features.
3.3 Feature Discrimination
Figure 4: Visual illustration of t-SNE clustering for image and point cloud features as learned by the model, shown (a) without and (b) with the incorporation of learnable transformation alignment.
Building dense point-to-pixel correspondences relies on learning discriminative features. Previous work [49] simply uses the cosine distance as descriptor loss to bring point and image features of positive pairs closer and those of negative pairs farther apart. The loss function drives the matching features to be similar. Although such a method tries to minimize the distribution gaps between the two different modalities by feature fusion, it is impossible to fully eliminate such discrepancy.
To align the two modalities, we apply an InfoNCE [50] loss on the fused features. Specifically, we sample a subset of points $\mathbf{P}_{S}$ inside the overlapped areas with the RGB image, then we can compute the InfoNCE loss for these points as:
$\mathcal{L}_{f}^{\mathbf{P}}=\,-\mathbb{E}_{\mathbf{p}_{i}\in\mathbf{P}_{S}}\left[\mathrm{log}\frac{f(\mathbf{p}_{i},\mathbf{I}_{\mathbf{p}_{i}})}{f(\mathbf{p}_{i},\mathbf{I}_{pr(i)})+\sum_{\mathbf{I}_{pr(i)}^{-}}f(\mathbf{p}_{i},\mathbf{I}_{pr(i)}^{-})}\right],$ (1)
where $pr(i)$ is the index of pixel corresponding to point $\mathbf{p}_{i}$, $\mathbf{I}_{pr(i)}$ and $\mathbf{I}_{pr(i)}^{-}$ denote pixels that match and do not match point $\mathbf{p}_{i}$, respectively. These pairs are determined according to the distance between the projection of points and pixels on the image plane. We denote the positive and negative margins as $r_{p}$ and $r_{n}$.
Here, we adopt the log-bilinear model for $f(\cdot,\cdot)$ following  [50]. To address the disparity in features between different modalities, we implement a learnable linear transformation (LTA) on the features from both modalities before computing the dot product. This approach is chosen over the traditional method of calculating feature similarities using cosine distance directly. The formula is expressed as:
$f(\mathbf{p}_{i},\mathbf{I}_{j})=e^{(\widetilde{\mathbf{F}}_{i}^{\mathbf{p}}\mathbf{W}_{f}\widetilde{\mathbf{F}}^{\mathbf{I}\top}_{j})/\tau}=e^{(\widetilde{\mathbf{F}}_{i}^{\mathbf{p}}\mathbf{Q}\mathbf{D}\mathbf{Q}^{\top}\widetilde{\mathbf{F}}^{\mathbf{I}\top}_{j})/\tau},$ (2)
where $\mathbf{W}_{f}$ represents a learnable linear transformation matrix, $\tau$ is the temperature factor, and $\mathbf{Q}\mathbf{D}\mathbf{Q}^{\top}$ is the decomposition result of $\mathbf{W}_{f}$, $\widetilde{\mathbf{F}}^{\mathbf{I}}$ and $\widetilde{\mathbf{F}}^{\mathbf{P}}$ denote normalized image and point features. In the implementation, we ensure the symmetry of $\mathbf{W}_{f}$ such that it can be decomposed according to the properties of spectral decomposition. And computing similarities between features, i.e., $\widetilde{\mathbf{F}}_{i}^{\mathbf{p}}\mathbf{W}_{f}\widetilde{\mathbf{F}}^{\mathbf{I}\top}_{j}$, is facilitated to be bidirectional. Similarly, we can obtain the discriminative loss $\mathcal{L}_{f}^{\mathbf{I}}$ for the image. This method, termed learnable transformation alignment, enables the alignment of features across different modalities through a learnable transformation.
We illustrate the learned features using cosine distance and learnable transformation alignment, respectively, in Fig. 4. We can observe that it is non-trivial to drive the matching 2D and 3D features to be similar with respect to cosine distance due to the inherent domain gap between point clouds and images. In contrast, when we use the learnable transformation alignment for computing similarities during training, the 2D and 3D features are closer and evenly distributed in the transformed unified representation space. We further analyze the effect of learnable transformation alignment in Section 4.5.
3.4 Building Dense Correspondence
Overlapping Area Detection. In our 2D-3D neural calibration framework, an essential component is overlapping area detection. Due to differences in the operational principles and the field of view of the camera and LiDAR sensors, the image and point cloud do not perfectly overlap. This module is responsible for identifying regions where the point cloud and image data overlap [49]. Accurate detection of these overlapping areas is crucial for effective feature alignment and calibration accuracy. Specifically, we detect the overlapping areas for input images and point clouds based on their corresponding fused features $\mathbf{F}^{\mathbf{I}}$ and $\mathbf{F}^{\mathbf{P}}$ [49]. We denote the $\mathbf{S}^{\mathbf{P}}$ and $\mathbf{S}^{\mathbf{I}}$ as the predicted probabilities of points and pixels in overlapping areas, respectively.
The points and pixels with scores higher than pre-set thresholds are considered inside the overlapping areas. We denote the estimated points and pixels in the overlapping areas as $\mathbf{P}_{o}\in\mathbb{R}^{N_{\mathbf{P}}\times 3}$ and $\mathbf{I}_{o}\in\mathbb{R}^{N_{\mathbf{I}}\times 3}$, where ${N_{\mathbf{P}}}$ and ${N_{\mathbf{I}}}$ are the number of points and pixels in the overlapping area.
During the training phase, the model supervises predicted overlap scores through binary cross-entropy loss. The loss function of overlapping area detection is defined as follows:
[…]
where $\mathbb{I}_{o}()$ is an indicator function that equals 1 if the point $\mathbf{p}_{i}$ or pixel $\mathbf{I}_{j}$ is in the overlapping area and 0 otherwise.
Soft Matching. Previous work simply applies the hard assignment strategy, i.e., non-differentiable argmax operation, to build correspondences between the points and pixels, i.e., assigning points in the overlapped area to the pixels most similar to them [49]. To make the overall framework end-to-end differentiable, here we predict the location of points in the image plane as a weighted sum of the target pixel coordinates:
$\widehat{\mathbf{C}}_{i}^{\mathbf{P}}=\sum_{\mathbf{I}_{j}\in\mathbf{I}_{o}}w_{ij}\cdot\mathbf{C}_{j}^{\mathbf{I}},$ (4)
where $\widehat{\mathbf{C}}_{i}^{\mathbf{P}}$ represents the predicted coordinate of the point cloud point $\mathbf{p}_{i}$ in the image plane, and $\mathbf{C}_{j}^{\mathbf{I}}$ denotes the 2D coordinate of the $j$-th pixel. The weight $w_{ij}$ signifies the feature similarity between the point $\mathbf{p}_{i}$ and the pixel $\mathbf{I}_{j}$, and is calculated based on the soft assignment principle:
$w_{ij}=\frac{\exp(\widetilde{\mathbf{F}}_{i}^{\mathbf{p}}\mathbf{W}_{f}\widetilde{\mathbf{F}}^{\mathbf{I}\top}_{j})}{\sum_{\mathbf{I}_{k}\in\mathbf{I}_{o}}\exp(\widetilde{\mathbf{F}}_{i}^{\mathbf{p}}\mathbf{W}_{f}\widetilde{\mathbf{F}}^{\mathbf{I}\top}_{k})}.$ (5)
This softmax-based weighting scheme ensures that the contribution of each pixel to the final predicted point location is proportional to its similarity with the point cloud point. Using the softmax function in this context allows for a differentiable and probabilistic approach to assigning correspondences, as opposed to the hard assignment methods used in previous works.
3.5 Estimation of Rigid Transformation
The final module in our framework estimates the rigid transformation based on the soft matching results computed in Section 3.4. The neural calibration problem is defined as determining the rigid transformation $\{\mathbf{R},\mathbf{t}\}$ that best maps points in overlapping areas $\mathbf{P}_{o}$ onto the estimated 2D coordinates in the image plane, i.e.,
$\widehat{\mathbf{R}},\widehat{\mathbf{t}}=\mathop{\arg\min}_{\mathbf{R},\mathbf{t}}\sum_{\mathbf{p}_{i}\in\mathbf{P}_{o}}\left[\mathbf{K}(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})-\widehat{\mathbf{C}}_{i}^{\mathbf{P}}\right].$ (6)
In this paper, we solve Eq. (6) using a differentiable EPnP algorithm [51].
The rigid transformation estimation is optimized by the loss function defined as:
$\mathcal{L}{p}={\left\|\mathbf{R}_{gt}^{\top}\widehat{\mathbf{R}}-\mathbf{E}\right\|}_{H}+{\left\|\mathbf{t}_{gt}-\widehat{\mathbf{t}}\right\|}_{H},$ (7)
where ${\left\|\cdot\right\|}_{H}$ is the Huber-loss, $\{\mathbf{R}_{gt},\mathbf{t}_{gt}\}$ indicates the ground-truth pose, $\mathbf{E}$ is an identify matrix.
3.6 Training Pipeline
Initially, each pair of an image and a point cloud is associated with a relative pose, denoted as $\{\mathbf{R}_{raw},\mathbf{t}_{raw}\}$. Since the camera and LiDAR sensors are fixed in a relatively close position of the cars, the original transformation from the camera coordinate to the point cloud coordinate is similar. To avoid the network overfitting to such scenarios, we simulate a range of sensor placements and orientations by applying random rotation $\mathbf{R}_{r}$ and translation $\mathbf{t}_{r}$ to the original point cloud during the pre-training process. The ground-truth pose is recalculated to accommodate these adjustments:
$\mathbf{t}_{gt}$ $=\mathbf{t}_{raw}-\mathbf{R}_{raw}\mathbf{t}_{r},$ (8)
$\mathbf{R}_{gt}$ $=\mathbf{R}_{raw}\mathbf{R}_{r}^{-1}.$
Our network is trained end-to-end, with supervision from the ground truth poses $\{\mathbf{R}_{gt},\mathbf{t}_{gt}\}$. The loss function is a weighted sum of three components: $\mathcal{L}=\lambda_{f}\mathcal{L}_{f}+\lambda_{o}\mathcal{L}_{o}+\lambda_{p}\mathcal{L}_{p}$. These components are weighted as follows: $\lambda_{f}=1.0$, $\lambda_{o}=0.5$, and $\lambda_{p}=0.2$. This training strategy ensures that our network learns to accurately estimate poses under varied conditions, enhancing the generalizability and effectiveness of learned representation.
4 Experiments
In this section, we begin with an overview of the datasets and evaluation metrics in Sec.4.1. Following this, Sec.4.2 to Sec.4.4 detail the experimental setup and the fine-tuning results for three downstream tasks. Subsequent sections, Sec.4.5 and Sec.4.6, delve into comprehensive ablation studies and in-depth analyses to assess critical aspects of our framework. Finally, we explore the reasons why the pre-text task is effective in Sec 4.8.
4.1 Dataset
In this part, we briefly describe the datasets used for pre-training and fine-tuning downstream tasks.
SemanticKITTI (SK) dataset contains RGB image and point cloud data pairs from KITTI scenes for the task of urban scene semantic segmentation [19]. The data is collected from sensor systems mounted on a car, comprising over 200,000 images and corresponding point clouds from 21 different scenes/sequences. The images and point clouds are synchronized to ensure a fixed relative transformation. The original images have a resolution of 1241x376 pixels. Each point cloud contains approximately 40,000 3D points. Following common practice, the dataset is split into a training set using the 10 sequences and a validation set using the eighth sequence.
TABLE I: Results of semantic segmentation models fine-tuned on three distinct datasets. This comparison considers the quantity of annotated data, the dataset employed for pre-training, and variations in model architecture. Our analysis contrasts NCLR with a baseline that lacks pre-training and other contemporary self-supervised methods, using the mean Intersection over Union (mIoU) percentage metric.
Dataset Backbone Method 0.1% 1% 10% 50% 100%
nuScenes [21] MinkUNet [37] No pre-training 21.6 35.0 57.3 69.0 71.2
PointContrast [3] 27.1 +5.5 37.0 +2.0 58.9 +1.6 69.4 +0.4 71.1 -0.1
DepthContrast [8] 21.7 +0.1 34.6 -0.4 57.4 +0.1 69.2 +0.2 71.2 0
ALSO [15] 26.2 +4.6 37.4 +2.4 59.0 +1.7 69.8 +0.8 71.8 +0.6
NCLR (Ours) 26.6 +5.0 37.8 +2.8 59.5 +2.2 71.2 +2.2 72.7 +1.5
SPVCNN [38] No pre-training 22.2 34.4 57.1 69.0 70.7
ALSO [15] 24.8 +2.6 37.4 +3.0 58.4 +1.3 69.5 +0.5 71.3 +0.6
NCLR (Ours) 25.8 +3.6 37.8 +3.4 59.2 +2.1 71.0 +2.0 73.0 +2.3
SemanticKITTI [19] MinkUNet [37] No pre-training 30.0 46.2 57.6 61.8 62.7
PointContrast [3] 32.4 +2.4 47.9 +1.7 59.7 +2.1 62.7 +0.9 63.4 +0.7
DepthContrast [8] 32.5 +2.5 49.0 +2.8 60.3 +2.7 62.9 +1.1 63.9 +1.2
SegContrast [5] 32.3 +2.3 48.9 +2.7 58.7 +1.1 62.1 +0.3 62.3 -0.4
ALSO [15] 35.0 +5.0 50.0 +3.8 60.5 +2.9 63.4 +1.6 63.6 +0.9
TARL [10] 37.9 +7.9 52.5 +6.3 61.2 +3.6 63.4 +1.6 63.7 +1.0
NCLR (Ours) 39.2 +9.2 53.4 +7.2 61.4 +3.8 63.5 +1.7 63.9 +1.2
SPVCNN [38] No pre-training 30.7 46.6 58.9 61.8 62.7
ALSO [15] 35.0 +4.3 49.1 +2.5 60.6 +1.7 63.6 +1.8 63.8 +1.1
NCLR (Ours) 38.8 +8.1 52.8 +6.2 61.1 +2.2 64.0 +2.2 64.1 +1.4
SemanticPOSS [22] MinkUNet [37] No pre-training 36.9 46.4 54.5 55.3 55.1
PointContrast [3] 39.3 +2.4 48.1 +1.7 55.1 +0.6 56.2 +0.9 56.2 +1.1
DepthContrast [8] 39.7 +2.8 48.5 +2.1 55.8 +1.3 56.0 +0.7 56.5 +1.4
SegContrast [5] 41.7 +4.8 49.4 +3.0 55.4 +0.9 56.2 +0.9 56.4 +1.3
ALSO [15] 40.7 +3.8 49.6 +3.2 55.8 +1.3 56.4 +1.1 56.7 +1.6
NCLR (Ours) 41.7 +4.8 49.8 +3.4 56.0 +1.5 56.6 +1.3 56.7 +1.6
KITTI 3D Object Detection Dataset (KITTI3D) is a prevailing collection of data specifically designed for 3D object detection in advancing autonomous driving technology [20]. The dataset is collected from various urban and suburban environments under different weather and lighting conditions. Each sample contains two modalities of 3D point clouds and RGB images. The LiDAR sensor used in the KITTI dataset is a Velodyne HDL-64E LiDAR. The FOV of the camera in the KITTI dataset aligns closely with the range of the LiDAR sensor, ensuring comprehensive coverage of the surroundings of the vehicle. Additionally, the dataset provides calibration information between the camera and the LiDAR sensor, which is essential for tasks that require the fusion of data from different sensors. The dataset includes several types of objects commonly encountered in driving scenarios, such as cars, pedestrians, and cyclists. Each object in the dataset is annotated with a 3D bounding box, which provides precise information about the object’s location, orientation, and dimensions in the 3D space. Following common practice, we further split all training samples into a subset with 3712 samples for training and the remaining 3769 samples for validation. We evaluate the performance by the Average Precision (AP) metric under IoU thresholds of 0.7, 0.5, and 0.5 for car, pedestrian, and cyclist categories, respectively. We computed APs with 40 sampling recall positions by default, instead of 11.
NuScenes Dataset is comprised of driving footage captured in Boston and Singapore using a vehicle outfitted with a 32-beam LiDAR sensor and other sensors [21]. Equipped with a full suite of sensors typical of autonomous vehicles, the dataset features a 32-beam LiDAR system, six cameras, and radar units, all providing a complete 360-degree coverage. The creators provide 850 total driving scene snippets, segmented into 700 training scene samples and 150 validation scene samples for the purposes of benchmarking models. Each of these scenes spans a duration of 20 seconds and is annotated at a frequency of 2Hz. The dataset provides detailed annotations for multiple object classes, such as vehicles, pedestrians, bicycles, and road barriers. Each object is annotated with a 3D bounding box, along with attribute information like visibility, activity, and pose. It has also been extended to include capabilities for semantic segmentation and panoptic segmentation, known as nuScenes-seg.
In an effort to enable fair evaluations against previously published models, we make use of 600 of the 700 training scenes to pretrain our model, reserving the leftover 100 training scenes to calibrate hyperparameters. After this tuning period, model performance is measured using the full 150 validation scenes provided in the dataset. This validation set therefore functions as our test set. It is useful to note that the nuScenes data includes LiDAR sweeps that have not been manually annotated - the sensors record at 20 Hz, while human annotations are only provided for every 10th LiDAR sweep. These unannotated scans could be utilized in a self-supervised manner to pretrain models. However, to stay consistent with prior work, we only leverage the annotated scans during our pre-training phase, while ignoring the annotations themselves. After pretraining, we fine-tune our model on various subsets of the 700 scene in the training set. More specifics on this precise evaluation protocol can be found in these published benchmark papers  [21]. In our experiments, we follow this same protocol to allow for standardized comparisons.
SemanticPOSS Dataset is a valuable asset for 3D semantic segmentation studies [22], consisting of 2988 diverse and complex LiDAR scans featuring numerous dynamic instances. Collected at Peking University, it conforms to the SemanticKITTI data format. This dataset is especially relevant for autonomous driving research, encompassing 14 categories like people, riders, and cars. In line with the official evaluation guidelines, the third sequence is designated as the validation set, while the remaining sequences form the training set.
Figure 5: Qualitative analysis of semantic segmentation on SemanticKITTI: fine-tuning performance with 1% of annotated training data. We highlight some areas that are easily misclassified with red circles.
4.2 Semantic Segmentation
Network Architectures. In order to assess the versatility of our method across various architectures and ensure an equitable comparison with prior studies, we conducted experiments using multiple backbone networks. Our tests included two versions of MinkUNet [37], along with ResUNet18 applied to SemanticKITTI, ResUNet34 for nuScenes, and the SPVCNN [38]. For the image branch, we utilize the ResNet backbone with the FPN structure to extract features. During the fine-tuning process for semantic segmentation, the image branch is omitted, and we substitute the final layer of the point cloud backbone. This replacement involves introducing a new fully-connected layer, designed with the channel matching the number of segmentation classes. This modification facilitates per-point predictions specifically adapted for the segmentation task at hand.
Pre-training Protocol. We pre-train the backbones for 50 epochs with a total batch size of 32 distributed across 8 GPUs for SemanticKITTI, and a batch size of 16 distributed on 4 GPUs for nuScenes dataset. We train the network with the AdamW optimizer [52], a learning rate of 0.001, a weight decay of 0.001, a temperature hyperparameter $\tau$ of 0.07 in Eq. (2), and a cosine annealing scheduler. During the pre-training process, all points, including those labeled as “ignore” are taken as input, and those not labeled points are filtered out for downstream fine-tuning. The input range on the x-y plane is [-51.2m, 51.2m] and [-3m, 1m] on the $z$-axis. Point clouds are voxelized with a grid size of 0.05m and 0.1m for SemanticKITTI and nuScenes datasets, respectively. During training, the image size is resized to 160x512 and 160x320 for SemanticKITTI and nuScenes, respectively. The point clouds undergo random transformations, including rotation around the $z$-axis and 2D translation in the $x$-$y$ plane, where the rotation angle and translation of each direction uniformly drawn from [-$\pi$, $\pi$] and [-15m, 15m], respectively.
Downstream Fine-tuning Protocol. Following the pre-training phase, we fine-tune the neural network backbones specifically for the task of semantic segmentation. This fine-tuning is conducted on selected data subsets from the SemanticKITTI and nuScenes datasets. We adhere to the fine-tuning hyperparameters and dataset splits as detailed in ALSO [15]. In the case of nuScenes, we implement a batch size of 8, and for SemanticKITTI, we opt for a smaller batch size of 2. The final per-point scoring is determined by assigning each point the prediction of the voxel in which it is located. It should be noted that the epoch count varies in relation to the percentage of training data used: for 0.1% of the data, 1000 epochs are used; for 1%, it’s 500 epochs; 100 epochs are applied to 10% of the data; 50 epochs for 50%; and 30 epochs when training with the full dataset.
Quantitative Evaluation Results. Table I presents the experimental results experiments on fine-tuning semantic segmentation models using three distinct datasets: nuScenes, SemanticKITTI, and SemanticPOSS. We observe that NCLR consistently outperformed other methods across different data volumes, showing particularly strong gains over the baseline training from scratch. For instance, with 0.1% of the data, NCLR achieved a 5.0% and 3.6% improvement in mIoU over the baseline for MinkUNet and SPVCNN, respectively. Similarly, NCLR demonstrated substantial improvements on the SemanticKITTI dataset, especially with limited training data. For MinkUNet, it showed a 9.2% increase in mIoU with just 0.1% of the data compared to the baseline. Similar trends were observed with SPVCNN. In the SemanticPOSS dataset, NCLR exhibits superior performance compared to traditional non-pre-trained models and other self-supervised approaches. The improvement is consistent across all data levels, demonstrating the generalization and effectiveness of NCLR in learning useful representations for 3D semantic segmentation tasks.
TABLE II: Comparisons between our method and other self-supervised learning methods fine-tuned on the KITTI3D dataset. We report the AP evaluated with 40 recall positions on the val set of the KITTI3D dataset. Note ProposalContrast [7] is specifically tailored for 3D object detection and pre-trained with the large-scale Waymo dataset.
Detector Pre-training Schedule Car Pedestrian Cyclist Avg
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
SECOND Scratch 90.20 81.50 78.61 53.89 48.82 44.56 82.59 65.72 62.99 67.65
ALSO [15] 90.20 81.53 78.83 57.30 53.21 48.32 82.92 69.12 64.57 69.56
NCLR (Ours) 90.23 81.99 79.05 59.20 54.75 49.32 83.64 70.16 65.13 70.38
PV-RCNN Scratch 91.74 84.60 82.29 65.51 57.49 52.71 91.37 71.51 66.98 73.80
ALSO [15] 92.15 84.85 82.59 65.63 57.83 53.14 91.81 73.85 69.71 74.62
STRL [11] - 84.70 - - 57.80 - - 71.88 - -
PointContrast [3] 91.40 84.18 82.25 65.73 57.74 52.46 91.47 72.72 67.95 73.99
ProposalContrast [7] 92.45 84.72 82.47 68.43 60.36 55.01 92.77 73.69 69.51 75.49
NCLR (Ours) 92.43 84.86 82.58 67.89 60.48 55.47 92.45 74.05 70.29 75.61
Qualitative Results. Fig. 5 illustrates qualitative results derived from various methods. Our self-supervised pre-training method produced more accurate segmentation results compared to other self-supervised learning approaches. The segmentation results from our method contain more precise boundaries and finer details. Specifically, it achieved better performance at distinguishing between visually similar object classes like the road and sidewalk.
TABLE III: The results (mAP) for models fine-tuned with varying quantities of annotated data. These results specifically reflect the performance in 3D object detection under moderate difficulty conditions on the validation set of the KITTI3D dataset.
Ratio Detector Pre-training Schedule Car Ped. Cyc. mAP (%)
10% SECOND Scratch 75.02 40.53 46.09 53.88
NCLR (Ours) 76.57 42.69 49.69 56.31(+2.43)
PV-RCNN Scratch 79.65 51.42 61.06 64.04
NCLR (Ours) 82.32 55.95 62.52 66.93(+2.89)
20% SECOND Scratch 78.12 42.35 60.97 60.48
NCLR (Ours) 79.17 44.62 64.78 62.85(+2.37)
PV-RCNN Scratch 82.37 53.70 67.31 67.79
NCLR (Ours) 82.56 57.29 69.92 69.92(+2.13)
50% SECOND Scratch 80.84 46.55 64.02 63.81
NCLR (Ours) 81.80 49.38 66.27 65.81(+2.00)
PV-RCNN Scratch 82.45 56.87 70.36 69.89
NCLR (Ours) 82.95 58.91 73.67 71.84(+1.95)
4.3 3D Object Detection
Network Architectures. In the downstream 3D object detection task, we explore two prevalent network architectures: the SECOND [44] and PV-RCNN [45] object detectors. Both architectures are built upon a shared foundational design, which includes a 3D sparse encoder. This encoder processes the input voxels through 3D sparse convolutional operations. In addition, they incorporate a bird’s-eye-view encoder (termed the 2D-backbone), which is activated post BEV projection. The key difference between them lies in their respective detection heads. SECOND employs a region proposal network (RPN) directly on the 2D backbone, whereas PV-RCNN enhances the RPN predictions through point-level refinement. This refinement leads to more accurately defined bounding boxes and improved confidence estimations.
Pre-training Protocol. In alignment with the requirements of downstream detection tasks, we process the point clouds through a voxelization step. This involves setting the grid size to [0.05m, 0.05m, 0.1m] along the $x,y$ and $z$ axes respectively. The maximum number of non-empty input voxels is limited to 60,000. The raw image is resized to (512, 160) as input. We utilize the default AdamW optimizer. The settings of the optimizer include a peak learning rate of 0.001 and a weight decay factor of 0.001. Cosine learning rate schedule [53] is adopted. We pretrain the backbone for 50 epochs on SemanticKITTI dataset with a batch size of 8 on a single GPU. Regarding the configuration of the VoxelNet backbone [43], we ensure consistency by employing parameters identical to those used in the corresponding 3D object detectors.
Downstream Fine-tuning Protocol. In the subsequent stage, the detection module from either SECOND or PVRCNN is integrated with the pre-trained neural network, and the combined detector is further fine-tuned on the KITTI3D dataset. This process utilizes the OpenPCDet framework [^1] https://github.com/open-mmlab/OpenPCDet, specifically its implementation of these detectors, along with the standard training parameters provided by OpenPCDet. Consistent with the methodology outlined in a prior study [15], this fine-tuning process is repeated three times independently, and the highest mean Average Precision (mAP) achieved on KITTI3D’s validation set is recorded and presented.
TABLE IV: The performance metrics (PQ and IoU) for panoptic segmentation on SemanticKITTI showcase the effectiveness of the pre-trained models when fine-tuning with various proportions of annotated data.
Pre-training Schedule 0.10%   1%   10% 50% 100%
PQ IoU PQ IoU PQ IoU PQ IoU PQ IoU
From Scratch 4.76 11.13 22.72 30.84 47.20 53.53 55.32 61.94 55.40 59.75
PointContrast [3] 5.86 11.51 27.37 32.49 47.57 54.63 54.21 59.48 55.85 61.49
DepthContrast [8] 7.65 13.56 27.31 32.30 46.85 51.27 54.55 59.60 56.15 60.81
SegContrast [5] 7.58 14.46 26.14 32.85 47.02 53.47 55.38 60.04 56.73 61.96
TARL [10] 10.26 17.01 29.24 34.71 51.27 57.59 56.10 62.36 56.57 62.05
NCLR (Ours) 12.79 19.14 30.69 36.24 52.79 58.78 56.78 62.54 56.87 62.69
TABLE V: Experimental results from models pre-trained through self-supervised and supervised training on SemanticKITTI, subsequently fine-tuned for panoptic segmentation tasks using both the mini and complete training sets of nuScenes. We report the PQ and IoU metrics on the full validation set of nuScenes.
Pre-training Schedule Mini Full
PQ IoU PQ IoU
From Scratch 23.78 23.96 52.98 58.17
Supervised pre-training 24.77 23.6 53.19 58.05
PointContrast [3] 26.58 25.46 51.06 56.39
DepthContrast [8] 28.66 27.3 51.51 57.06
SegContrast [5] 28.84 26.79 52.31 57.24
TARL [10] 32.22 30.73 53.26 59.14
NCLR (Ours) 33.37 31.62 53.97 59.76
Quantitative Evaluation Results. We evaluate the transferability of our pre-trained model by first pre-training on the SemanticKITTI dataset and subsequently fine-tuning on the KITTI3D dataset. We report the experimental results in Table II. It is evident that our approach consistently surpasses the baseline established by training from scratch, achieving an enhancement for both SECOND and PV-RCNN. With our pre-training method, the performance improves, particularly in the Pedestrian and Cyclist categories, reaching an average mAP of 70.38%. Our method performs comparably to state-of-the-art models like ProposalContrast [7], which is specifically tailored for detection tasks and pre-trained with the large-scale Waymo dataset.
One of the primary benefits of self-supervised learning lies in its ability to improve the performance of downstream tasks when the annotation source is limited. To further assess this aspect, we evaluated our method in label-efficient 3D object detection. We consider a model trained from a state of random initialization as our standard for comparison. The superiority of our pre-trained model compared to this standard is detailed in Table III. Essentially, our pre-trained model consistently boosts the performance in detection tasks using both SECOND and PV-RCNN architectures, particularly notable under conditions of limited labeled data – showing an improvement of 2.43% and 2.89% in mAP with just 1% labeled data for SECOND and PV-RCNN, respectively. Besides, our model surpasses the baseline performances across all variations of available annotated data quantities.
4.4 Panoptic Segmentation
To assess the instance-level features learned by our method, we further fine-tune our pre-trained models for the task of panoptic segmentation.
Network Architectures. In this part, we choose the MINKUnet-34 [37] as the 3D point-wise feature backbone. For the fine-tuning process, semantic and instance heads are integrated with the pre-trained 3D backbone, followed by clustering post-processing to identify the instances [10]. This clustering stage leverages semantic predictions to exclude background points, focusing solely on foreground instances such as vehicles, pedestrians, and cyclists. Once the irrelevant points are removed, the remaining points undergo a clustering process to discern distinct instances based on the features from the instance head. For the clustering phase, we opt for the mean shift algorithm  [54] with a set bandwidth of 1.2 and a minimum cluster size requirement of 50 points. The other architectures and settings are the same as Section 4.2 during pre-training.
Protocol for Fine-tuning on Downstream Tasks. Our methodology employs the AdamW optimizer, utilizing a max learning rate set at 0.2. In the context of panoptic segmentation applied to the SemanticKITTI dataset, the fine-tuning of our model is conducted across various subsets of the annotated training data, specifically 0.1%, 1%, 10%, 50%, and the full 100% subset. For the nuScenes dataset, we fine-tune using both the complete training set and the mini training subset that is provided. The performance of the fine-tuned model is then assessed on the complete validation set of all datasets. We fine-tune the model for fixed 50 epochs.
Results of Fine-tuning. We pre-train the backbone on the SemanticKITTI dataset and fine-tune it on different percentage subsets of SemanticKITTI dataset. As shown in Table IV, our method is consistently better than previous self-supervised pre-training approaches. When the segmentation model is trained with fewer labels, we can observe an obvious improvement compared to the baseline of scratch training.
Generalization of Learned Features. Our study also investigates the adaptability of learned features focusing on panoptic segmentation tasks. We initially pre-train the 3D backbone on the SemantiKITTI dataset and subsequently fine-tune it on both the full and mini training sets of nuScenes. As presented in Table V, while all approaches enhance performance on the nuScenes dataset, the features generated through our technique demonstrate greater adaptability when employed in a different domain dataset. Furthermore, in comparison to supervised pre-training on SemanticKITTI, our approach exhibits improved results. This finding underscores the enhanced effectiveness and potential of our method in scenarios of transfer learning, surpassing the conventional supervised pre-training methods.
In general, the experiments conducted further confirm that our approach effectively extracts semantic nuances and identifies instance-level details, achieving higher scores in IoU and PQ metrics than former methodologies.
Figure 6: Visual comparison of the similarities between image and point cloud features learned by the model w/o and w/ learnable transformation alignment. For a selected 3D query point (indicated by a purple dot), we calculate its similarities with image features, forming a 2D similarity map. These similarity maps are illustrated using examples from the SemanticKITTI validation set.
Figure 7: Visual comparison of the similarities between image and point cloud features learned by the model w/o and w/ learnable transformation alignment. For a selected 2D query pixel (marked with a purple dot), we calculate its similarities with point features and project these similarities onto the image plane, forming a similarity map. These similarity maps are illustrated using examples from the SemanticKITTI validation set.
4.5 Ablation Studies
In our ablation studies on the nuScenes dataset, we adhere to the evaluation protocol from ALSO [15]. This involves splitting the nuScenes training set into two parts: ablation-train and ablation-val. Fine-tuning is performed with just 1% of annotations from the ablation-train set. This ensures parameter tuning does not involve the validation set, reserved for comparing against other methods. We limit the training period for these studies to 100 epochs.
TABLE VI: Ablation study on nuScenes semantic segmentation. End-to-end pose estimation indicates using soft-matching, differentiable EPnP solver, and loss of pose estimation.
Baseline
Learnable Transformation Alignment $\times$
Overlapping Area Detection $\times$ $\times$
End-to-End Pose Estimation $\times$ $\times$ $\times$
mIoU (%) 35.90 36.52 37.13 37.92
Effect of Key Components. We analyze the impact of each component incorporated into our method in Table VI,. We can observe that our proposed key components benefit the pre-training framework in learning useful representation and yield better performance in downstream tasks. For example, the introduction of learnable transformation alignment increases the mIoU from 35.90% to 36.52%. This indicates that allowing the model to adaptively transform features enhances its ability to capture the inherently non-linear and complex relationship between point cloud data and RGB images. Besides, the supervision of advanced tasks, including overlapping area detection and pose estimation also prompts the network to fully understand the two modalities and learn useful representation. Then we analyze the effect of learnable transformation alignment in detail.
TABLE VII: Comparison of similarities of image and corresponding point features learned w/o and w/ learnable transformation alignment (LTA). We measure the mean and standard value of similarities on the validation set of SemanticKITTI.
Method Similarity
w/o LTA 0.55 $\pm$ 0.17
w/ LTA 0.69 $\pm$ 0.13
Figure 8: Evaluating the accuracy of feature matching across various techniques. (a) Similarities between image and point features are determined using cosine distance, without the aid of learnable transformation alignment, and correspondences between points and pixels are established through hard assignment. (b) In contrast, this approach incorporates learnable transformation alignment for calculating feature similarities and employs soft-matching to establish the relationships between points and pixels. The matching accuracy reflects the discrepancy between the locations of matched pixels and the actual projections of points on the image plane.
Effect of Learnable Transformation Alignment. The gap between the features of the two modalities – point clouds and RGB images – is a significant challenge in computing their similarity for matching purposes (see Fig. 4). By transforming features from both modalities into a common feature space, the learnable transformation $\mathbf{W}_{f}$ helps in aligning these features more effectively. We further visualize the similarities between image and point features in Fig. 6 and Fig. 7.
We also provide the quantitative comparison of similarities between features of two modalities in Table VII. Specifically, we compute the cosine similarity between the features of the points and corresponding pixels with and without linear feature transformation, respectively. We can observe that directly computing the cosine similarities yields smaller values for the corresponding points and pixels. However, applying the learnable feature transformation bridges the gap between the two modalities and brings more reasonable similarities.
TABLE VIII: Feature matching accuracy on SemanticKITTI dataset. The error indicates the discrepancy between the locations of matched pixels and the actual projections of points on the image plane. “Cosine” denotes directly computing the cosine distance of image and point features as similarities. “Learnable” denotes learnable transformation alignment. “Acc.” indicates the percentage of fine matchings with errors less than 5 pixels.
Method Error (pixel) Acc. (%)
Cosine + Hard-Match 4.58 $\pm$ 4.70 69.22
Cosine + Soft-Match 3.91 $\pm$ 4.06 74.27
Learnable + Hard-Match 3.08 $\pm$ 2.40 83.18
Learnable + Soft-Match 2.34 $\pm$ 1.05 90.28
4.6 Feature Matching Accuracy
We present a comparative analysis of feature matching accuracy using various feature alignment and point-to-pixel matching methods. The mean and standard deviation of error distances are detailed in Table VIII. To assess the precision of different methods, we calculate the matching accuracy (Acc.), defined as the percentage of fine matchings with errors less than 5 pixels. Results in Table VIII demonstrate that the integration of learnable transformation alignment and the soft-matching approach enhances matching accuracy on the SemanticKITTI dataset. Consequently, our method facilitates the network’s comprehensive understanding of the relationship between image and point cloud data, learning effective representation to process information from each domain.
Additional visualization of feature matching is presented in Fig. 8. Within the overlapping region, we determine the corresponding 2D pixel for each point and evaluate the error. This is done by initially projecting the point onto the image space and then calculating the Euclidean distance between the projected point and its matched 2D pixel. It is evident from our observations that the implementation of our proposed learnable transformation alignment combined with soft-matching markedly reduces the feature-matching error, in comparison to approaches that do not utilize these methods. Notably, in most instances, the errors are confined to within a 3-pixel range.
Figure 9: Illustration of the registration recall performance of various methods against different Relative Translation Error (RTE) and Relative Rotation Error (RRE) thresholds on the SemanticKITTI datasets. The $x$-axis represents the threshold levels for RTE and RRE, while the $y$-axis indicates the recall rate, reflecting the proportion of samples where RREs or RTEs remain below the corresponding threshold.
Figure 10: Visual results of detecting overlapping areas between images and point clouds. The first row depicts the actual projection of the point cloud onto the image plane, with varying colors indicating the depth of the points. The detection results on the image and point cloud are shown in the second and third rows, respectively. In these rows, green pixels and points represent correct detections within the overlapping regions of the image and point cloud. Conversely, red indicates the pixels and points are incorrectly classified in the overlapping areas, and yellow highlights areas are incorrectly identified as non-overlapping.
Figure 11: The visual results from the 2D-3D neural calibration on the SemanticKITTI dataset. In the first columns, we show the serious misalignment between the image and point cloud, which has initially undergone random rotation and translation. In the second column, we project the point cloud on the image plane with our estimated spatial transformation between the camera and LiDAR systems. Meanwhile, the third column illustrates the projections of the point cloud with ground-truth rigid transformation (camera pose). The painted color indicates the depth of each LiDAR point.
4.7 Accuracy of Image-to-Point Cloud Calibration
We compared our approach with other leading methods for image-to-point cloud registration/calibration [55, 49, 56]. Following standard practices in this field, we used the SemanticKITTI dataset to measure the accuracy of our method, employing metrics such as Relative Translational Error (RTE) and Relative Rotational Error (RRE) [49]. We provided a detailed analysis by illustrating the registration recall at various RTE and RRE thresholds on the SemanticKITTI dataset, as shown in Fig. 9. The results clearly indicate the enhanced performance of our technique in terms of registration accuracy.
While numerous studies have addressed the task of image-to-point cloud registration, it is noteworthy that these methods predominantly utilize simulated data. In these simulations, point clouds are subjected to random rotation and translation transformations. Consequently, the transformed point clouds are not aligned with the corresponding images. However, the transformations applied in these simulations are idealized, restricted to limited axes and not reflective of the general camera poses found in realistic scenes. This limitation potentially reduces the practical applicability of such methods in real-world scenarios, such as robot localization. In our research, we broaden the scope of image-to-point cloud registration/calibration by extending its application to self-supervised pre-training for 3D perception tasks.
4.8 Further Discussions
Fig. 10 demonstrates the precise detection of overlapping areas in both image and point cloud domains. In the image domain, inaccuracies are mostly confined to boundary areas. On the side, Fig. 11 illustrates the registration performance on the SemanticKITTI dataset. The figures in first column reveal noticeable misalignments between the image and the point cloud initialized with random rotation and translation. Despite the challenging initial conditions, our method accurately estimates the applied rigid transformation. This precise detection of overlapping areas and accurate estimation of rigid transformations indicate that the network could effectively understand the information and relationships between each domain. Such insights help explain why the representations learned in the pretext task are advantageous for downstream tasks.
5 Conclusion
We have introduced an innovative perspective to self-supervised learning by achieving a thorough alignment between two distinct modalities. In the pretext task of 2D-3D neural calibration, this network not only learns fine-grained matching from individual points to pixels but also achieves a comprehensive alignment between the image and point cloud data, i.e., understanding their relative pose. The integration of a soft-matching strategy alongside a differentiable PnP solver makes our framework end-to-end differentiable, thereby facilitating a more comprehensive understanding of multi-modal data. Besides, our method overcomes the limitations of traditional contrastive learning by introducing a transformation alignment technique that effectively mitigates the domain gap between image and point cloud data. The efficacy of our proposed pre-training method is substantiated through extensive experiments across various datasets and fine-tuning tasks, including LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation. Our results indicate a notable superiority over existing methods, underscoring the potential of our approach in enhancing 3D perception tasks.

### v2

@@
 This paper introduces a novel self-supervised learning framework for enhancing 3D perception in autonomous driving scenes.
-Specifically, our approach, named NCLR, focuses on 2D-3D neural calibration, a novel pretext task that estimates the rigid transformation aligning camera and LiDAR coordinate systems.
+Specifically, our approach, namely NCLR, focuses on 2D-3D neural calibration, a novel pretext task that estimates the rigid pose aligning camera and LiDAR coordinate systems.
 First, we propose the learnable transformation alignment to bridge the domain gap between image and point cloud data, converting features into a unified representation space for effective comparison and matching.
 Second, we identify the overlapping area between the image and point cloud with the fused features.
-Third, we establish dense 2D-3D correspondences to estimate the rigid transformation.
+Third, we establish dense 2D-3D correspondences to estimate the rigid pose.
 The framework not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, understanding their relative pose.
-We demonstrate NCLR’s efficacy by applying the pre-trained backbone to downstream tasks, such as LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation.
+We demonstrate the efficacy of NCLR by applying the pre-trained backbone to downstream tasks, such as LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation.
 Comprehensive experiments on various datasets illustrate the superiority of NCLR over existing self-supervised methods.
 The results confirm that joint learning from different modalities significantly enhances the network’s understanding abilities and effectiveness of learned representation.
-Code will be available at https://github.com/Eaphan/NCLR.
+The code is publicly available at https://github.com/Eaphan/NCLR.
 Index Terms:
-Self-supervised Learning, 3D Perception, Cross-modal, Autonomous Driving.
-1 Introduction
+Self-supervised Learning, 3D Perception, Cross-modal, Autonomous Driving, Registration.
+I Introduction
 The LiDAR technology serves as a vital enhancement to 2D cameras by precisely capturing the surroundings of a vehicle, offering robust performance in challenging conditions, including low light, intense sunlight, or glare from approaching headlights.
 This 3D perception, derived from LiDAR point clouds, is essential for the effective navigation of autonomous vehicles in three-dimensional spaces.
-While current leading methods are based on extensive labeled datasets, the process of labeling 3D data is exceedingly expensive and time-intensive, given the limited availability of annotation resources.
+While current leading methods are based on extensive labeled datasets, labeling 3D data is exceedingly expensive and time-intensive, given the limited availability of annotation resources.
 Consequently, there is a growing imperative to utilize unlabeled data.
@@
 Figure 1: Illustration of the proposed 2D-3D neural calibration developed as the pretext task for self-supervised learning of 3D LiDAR point clouds.
-Our method not only learns the local matching from points to pixels via contrastive losses, but also estimates the (unknown) holistic rigid transformation between camera and LiDAR systems.
+Our method not only learns the local matching from points to pixels via contrastive losses but also estimates the (unknown) holistic rigid pose aligning the camera and LiDAR systems.
 Besides, we propose learnable transformation alignment to fill the domain gap between the image and the point cloud.
@@
 During pre-training, our pretext task not only learns fine-grained matching from individual points to pixels but also achieves a comprehensive alignment between the image and point cloud data, i.e., understanding their relative pose.
-Specifically, to impose supervision on the rigid transformation estimation, we propose an end-to-end differentiable framework distinguished by its integration of a soft-matching strategy and a differentiable PnP solver.
+Specifically, to impose supervision on the rigid pose estimation, we propose an end-to-end differentiable framework distinguished by its integration of a soft-matching strategy and a differentiable PnP solver.
 We posit that this joint learning from different modalities for both local and global-level alignment will enhance the network’s capacity for sophisticated understanding and enable it to develop effective representations.
@@
 The rest of this paper is organized as follows.
-Section 2 provides an overview of related literature pertinent to our research.
-In Section 3, we present the overall architecture of NCLR and elaborate its principal elements.
-Section 4 presents an empirical evaluation of our proposed approach across three distinct downstream tasks, along with ablation studies to assess the impact of key components.
-The paper concludes with Section 5, summarizing our findings.
-2 Related Work
-2.1 Self-supervised Learning for Images
+Section II provides an overview of related literature pertinent to our research.
+In Section III, we present the overall architecture of NCLR and elaborate its principal elements.
+Section IV presents an empirical evaluation of our proposed approach across three distinct downstream tasks, along with ablation studies to assess the impact of key components.
+The paper concludes with Section V, summarizing our findings.
+II Related Work
+II-A Self-supervised Learning for Images
 Self-supervised learning involves training models on unlabeled data by generating pseudo-labels or tasks based on the data itself, without human annotations.
@@
 The field of self-supervised learning on images has seen significant evolution over the years.
-Initial approaches were centered around pretext tasks, which are auxiliary tasks designed to derive meaningful data representations.
-These tasks involved teaching models to restore color in black-and-white images [23] and to ascertain the spatial relation between image segments [24].
-The advent of deep learning brought about a paradigm shift in the field.
+Initial approaches are centered around pretext tasks, which are auxiliary tasks designed to derive meaningful data representations.
+These tasks involve teaching models to restore color in black-and-white images [23] and to ascertain the spatial relation between image segments [24].
+The advent of deep learning has brought about a paradigm shift in the field.
 One such approach is contrastive learning [25, 26].
@@
 Generative models [2, 27] have also emerged as a significant trend in self-supervised learning.
-These models, often employing Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), learned to generate images that closely resemble the training data.
-By learning to generate images, these models captured the underlying data distribution, which can be leveraged for various downstream tasks.
+These models, often employing Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), learn to generate images that closely resemble the training data.
+By learning to generate images, these models capture the underlying data distribution, which can be leveraged for various downstream tasks.
 In addition to these, there have been several exciting advancements in the field.
@@
 Multi-modal models aim to learn representations across different modalities (e.g., text, image, audio) using self-supervised learning.
-The Vision Transformer (ViT) adapted the transformer architecture, initially developed for natural language processing, to computer vision tasks [29].
+The Vision Transformer (ViT) adapts the transformer architecture, initially developed for natural language processing, to computer vision tasks [29].
 Additionally, there has been explored into incorporating physics-based priors in self-supervised learning.
-2.2 Self-supervised Learning for Point Clouds
+II-B Self-supervised Learning for Point Clouds
 The advances in self-supervision on point clouds have closely followed the improvements made in images.
-Early self-supervised methods used pretext tasks such as predicting transformations applied to the point cloud or reconstructing parts of the point cloud [30, 31].
-These methods were applied on dense scans of single objects [32, 33].
+Early self-supervised methods use pretext tasks such as predicting transformations applied to the point cloud or reconstructing parts of the point cloud [30, 31].
+These methods are applied on dense scans of single objects [32, 33].
 Discriminative-based Methods.
-This type of method explored different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduced contrastive losses to capture the geometric and structural nuances of 3D data [8, 9].
+This type of method explores different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduces contrastive losses to capture the geometric and structural nuances of 3D data [8, 9].
 Nunes et al.
@@
 Temporal-consistency-based Methods.
-STRL [11] utilized pairs of temporally-linked frames from 3D point cloud sequences, applying spatial data augmentation to learn invariant representations.
-STSSL [12] incorporated spatial-temporal positive pairs, introducing a point-to-cluster technique for spatial object distinction and a cluster-to-cluster method using unsupervised tracking for temporal correlations.
-TARL [10] used vehicle motion to align objects over time in different scans, focusing on enhancing segment-level feature similarities for consistent temporal representations.
-These works demonstrated how leveraging temporal information can lead to robust and invariant representations, which are beneficial for various downstream tasks.
+STRL [11] utilizes pairs of temporally-linked frames from 3D point cloud sequences, applying spatial data augmentation to learn invariant representations.
+STSSL [12] incorporates spatial-temporal positive pairs, introducing a point-to-cluster technique for spatial object distinction and a cluster-to-cluster method using unsupervised tracking for temporal correlations.
+TARL [10] uses vehicle motion to align objects over time in different scans, focusing on enhancing segment-level feature similarities for consistent temporal representations.
+These works demonstrate how leveraging temporal information can lead to robust and invariant representations, which are beneficial for various downstream tasks.
 Reconstruction-based Methods.
 Reconstruction-based methods have also been successful for self-supervision on point clouds.
-Some methods reconstructed point coordinates using the Chamfer distance [13, 14].
-Point-BERT randomly masked some patches and fed them into Transformers to recover the original tokens at the masked locations [34].
-Recently, ALSO [15] proposed to use unsupervised surface reconstruction as a pretext task to train 3D backbones on automotive LiDAR point clouds.
-Using the knowledge of occupancy before and after an observed 3D point along a LiDAR ray, it learned to construct an implicit occupancy function and good point features.
+Some methods reconstruct point coordinates using the Chamfer distance [13, 14].
+Point-BERT randomly masks some patches and feeds them into Transformers to recover the original tokens at the masked locations [34].
+Recently, ALSO [15] proposes to use unsupervised surface reconstruction as a pretext task to train 3D backbones on automotive LiDAR point clouds.
+Using the knowledge of occupancy before and after an observed 3D point along a LiDAR ray, it learns to construct an implicit occupancy function and good point features.
+Besides, rendering-based techniques [35, 36, 37] reconstruct masked scenes from point clouds or multi-view images, which have achieved promising generalization ability.
 Figure 3: The overall pipeline for 2D-3D neural calibration serves as a pretext task for generating self-supervised representation for LiDAR 3D point cloud data.
@@
 (4) Application of a soft-matching strategy to establish 2D-3D correspondences, followed by the use of a differentiable EPnP solver for camera pose estimation.
-Multi-modal Self-supervised Learning Another line of work leveraged synchronized and calibrated cameras and LiDAR to pretrain a 3D backbone.
-The underlying idea is to find pairs of corresponding points and pixels and ensure that the associated point and pixel representations are as close as possible [16, 17, 18].
-2.3 LiDAR-based 3D Perception
+Multi-modal Self-supervised Learning Another line of work leverages synchronized and calibrated cameras and LiDAR to pretrain a 3D backbone.
+These contrastive methods find pairs of corresponding points and pixels and ensure that the associated point and pixel representations are as close as possible [16, 17, 18].
+The primary motivation of these methods is to transfer the knowledge already embedded in a pre-trained 2D backbone to the 3D backbone.
+In contrast, our approach trains the network to achieve holistic alignment between the image and point cloud, enabling it to understand the two modalities thoroughly and subsequently learn effective 3D representations.
+Besides, our method does not rely on a pre-trained 2D backbone.
+By eliminating this dependency, our approach introduces greater flexibility and removes the constraint of inheriting specific representations from a 2D network.
+II-C LiDAR-based 3D Perception
 LiDAR-based 3D perception is a critical component in many applications, particularly in autonomous driving.
@@
 3D Semantic Segmentation.
-One of the foundational works in this area is PointNet [36], which marked a significant departure from previous methods by directly processing point clouds for tasks like segmentation.
-MinkUNet adapted the U-Net architecture for sparse 3D point clouds, using Minkowski Engine for efficient segmentation [37].
-SPVCNN combined sparse and continuous convolutions for scalable and efficient 3D point cloud segmentation [38].
+One of the foundational works in this area is PointNet [39], which marks a significant departure from previous methods by directly processing point clouds for tasks like segmentation.
+MinkUNet adapts the U-Net architecture for sparse 3D point clouds, using Minkowski Engine for efficient segmentation [40].
+SPVCNN combines sparse and continuous convolutions for scalable and efficient 3D point cloud segmentation [41].
 3D Object Detection.
-Advancements in 3D object detection have led to its categorization into voxel-based, point-based, and hybrid methods [39, 40, 41, 42].
+Advancements in 3D object detection have led to its categorization into voxel-based, point-based, and hybrid methods [42].
 Voxel-based techniques convert point clouds into uniform grid structures for feature extraction using sparse convolutions [43, 44].
@@
 introduced the Dynamic Shifting Network (DS-Net) as a potent framework for panoptic segmentation in the field of point clouds [47].
-This approach utilized cylinder convolution tailored for LiDAR data.
-The features derived were concurrently used by both the semantic and instance branches, functioning through a bottom-up clustering approach.
-3 Proposed Method
-3.1 Overview
+This approach utilizes cylinder convolution tailored for LiDAR data.
+The features derived are concurrently used by both the semantic and instance branches, functioning through a bottom-up clustering approach.
+III Proposed Method
+III-A Overview
 We propose a novel approach to generating self-supervised features for 3D point clouds, utilizing the 2D-3D neural calibration as a pretext task.
-The 2D-3D neural calibration aims to identify the unknown rigid transformation, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, through a neural network.
-This transformation aligns the coordinate systems of the camera and the point cloud.
+The 2D-3D neural calibration aims to identify the unknown rigid pose, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, through a neural network.
+This pose aligns the coordinate systems of the camera and the point cloud.
 We denote the input image as $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ and point cloud as $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}|i=1,2,...,N\}$, where $H$ and $W$ are the height and width of the image, and $N$ is the number of points.
@@
 The proposed framework is designed to be end-to-end trainable.
-3.2 Feature Extraction
+III-B Feature Extraction
 Both the image and point cloud branches employ an encoder-decoder structure.
@@
 If we need to pre-train a voxel-based backbone for downstream tasks, we can incorporate it and obtain the final point features by aggregating the voxel-wise features and fine-grained point-wise features.
-3.3 Feature Discrimination
+III-C Feature Discrimination
 Figure 4: Visual illustration of t-SNE clustering for image and point cloud features as learned by the model, shown (a) without and (b) with the incorporation of learnable transformation alignment.
@@
 In contrast, when we use the learnable transformation alignment for computing similarities during training, the 2D and 3D features are closer and evenly distributed in the transformed unified representation space.
-We further analyze the effect of learnable transformation alignment in Section 4.5.
-3.4 Building Dense Correspondence
+We further analyze the effect of learnable transformation alignment in Section IV-E.
+III-D Building Dense Correspondence
 Overlapping Area Detection.
@@
 Using the softmax function in this context allows for a differentiable and probabilistic approach to assigning correspondences, as opposed to the hard assignment methods used in previous works.
-3.5 Estimation of Rigid Transformation
-The final module in our framework estimates the rigid transformation based on the soft matching results computed in Section 3.4.
-The neural calibration problem is defined as determining the rigid transformation $\{\mathbf{R},\mathbf{t}\}$ that best maps points in overlapping areas $\mathbf{P}_{o}$ onto the estimated 2D coordinates in the image plane, i.e.,
-$\widehat{\mathbf{R}},\widehat{\mathbf{t}}=\mathop{\arg\min}_{\mathbf{R},\mathbf{t}}\sum_{\mathbf{p}_{i}\in\mathbf{P}_{o}}\left[\mathbf{K}(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})-\widehat{\mathbf{C}}_{i}^{\mathbf{P}}\right].$ (6)
-In this paper, we solve Eq. (6) using a differentiable EPnP algorithm [51].
-The rigid transformation estimation is optimized by the loss function defined as:
+III-E Estimation of Rigid Pose
+The final module in our framework estimates the rigid pose based on the soft matching results computed in Section III-D.
+The neural calibration problem is defined as determining the rigid pose $\{\mathbf{R},\mathbf{t}\}$ that best maps points in overlapping areas $\mathbf{P}_{o}$ onto the estimated 2D coordinates in the image plane, i.e.,
+$\widehat{\mathbf{R}},\widehat{\mathbf{t}}=\mathop{\arg\min}_{\mathbf{R},\mathbf{t}}\sum_{\mathbf{p}_{i}\in\mathbf{P}_{o}}\left[\mathbf{K}(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})-\widehat{\mathbf{C}}_{i}^{\mathbf{P}}\right],$ (6)
+where $\mathbf{K}$ denotes the camera intrinsic matrix.
+In this paper, we solve Eq. (6) using a differentiable EPnP algorithm [51]. The rigid pose estimation is optimized by the loss function defined as:
 $\mathcal{L}{p}={\left\|\mathbf{R}_{gt}^{\top}\widehat{\mathbf{R}}-\mathbf{E}\right\|}_{H}+{\left\|\mathbf{t}_{gt}-\widehat{\mathbf{t}}\right\|}_{H},$ (7)
 where ${\left\|\cdot\right\|}_{H}$ is the Huber-loss, $\{\mathbf{R}_{gt},\mathbf{t}_{gt}\}$ indicates the ground-truth pose, $\mathbf{E}$ is an identify matrix.
-3.6 Training Pipeline
+III-F Training Pipeline
 Initially, each pair of an image and a point cloud is associated with a relative pose, denoted as $\{\mathbf{R}_{raw},\mathbf{t}_{raw}\}$.
-Since the camera and LiDAR sensors are fixed in a relatively close position of the cars, the original transformation from the camera coordinate to the point cloud coordinate is similar.
+Since the camera and LiDAR sensors are fixed in a relatively close position of the cars, the original pose aligning the camera coordinate to the point cloud coordinate is similar.
 To avoid the network overfitting to such scenarios, we simulate a range of sensor placements and orientations by applying random rotation $\mathbf{R}_{r}$ and translation $\mathbf{t}_{r}$ to the original point cloud during the pre-training process.
@@
 This training strategy ensures that our network learns to accurately estimate poses under varied conditions, enhancing the generalizability and effectiveness of learned representation.
-4 Experiments
-In this section, we begin with an overview of the datasets and evaluation metrics in Sec.4.1.
-Following this, Sec.4.2 to Sec.4.4 detail the experimental setup and the fine-tuning results for three downstream tasks.
-Subsequent sections, Sec.4.5 and Sec.4.6, delve into comprehensive ablation studies and in-depth analyses to assess critical aspects of our framework.
-Finally, we explore the reasons why the pre-text task is effective in Sec 4.8.
-4.1 Dataset
+IV Experiments
+In this section, we begin with an overview of the datasets and evaluation metrics in Sec.IV-A.
+Following this, Sec.IV-B to Sec.IV-D detail the experimental setup and the fine-tuning results for three downstream tasks.
+Subsequent sections, Sec.IV-E and Sec.IV-F, delve into comprehensive ablation studies and in-depth analyses to assess critical aspects of our framework.
+Finally, we explore the reasons why the pre-text task is effective in Sec IV-H.
+IV-A Dataset
 In this part, we briefly describe the datasets used for pre-training and fine-tuning downstream tasks.
@@
 The data is collected from sensor systems mounted on a car, comprising over 200,000 images and corresponding point clouds from 21 different scenes/sequences.
-The images and point clouds are synchronized to ensure a fixed relative transformation.
+The images and point clouds are synchronized to ensure a fixed relative pose.
 The original images have a resolution of 1241x376 pixels.
@@
 Our analysis contrasts NCLR with a baseline that lacks pre-training and other contemporary self-supervised methods, using the mean Intersection over Union (mIoU) percentage metric.
+The best results are highlighted in bold.
 Dataset Backbone Method 0.1% 1% 10% 50% 100%
@@
 We highlight some areas that are easily misclassified with red circles.
-4.2 Semantic Segmentation
+TABLE II: Comparison of various pre-training techniques for semantic segmentation tasks using either finetuning or linear probing.
+We report the mean Intersection over Union (mIoU) on the nuScenes validation set across different annotation levels.
+Initialization Reference nuScenes SemanticKITTI
+Lin. Prob Finetune Finetune
+100% 1% 1%
+Random - 8.10 30.30 39.50
+PPKT [16] Arxiv’21 35.90 37.80 44.00
+SLidR [17] CVPR’22 38.80 38.30 44.60
+ST-SLidR [18] CVPR’23 40.48 40.75 44.72
+Ours - 42.81 42.29 45.84
+IV-B Semantic Segmentation
 Network Architectures.
@@
 The improvement is consistent across all data levels, demonstrating the generalization and effectiveness of NCLR in learning useful representations for 3D semantic segmentation tasks.
-TABLE II: Comparisons between our method and other self-supervised learning methods fine-tuned on the KITTI3D dataset.
+Qualitative Results.
+Fig.
+5 illustrates qualitative results derived from various methods.
+Our self-supervised pre-training method produced more accurate segmentation results compared to other self-supervised learning approaches.
+The segmentation results from our method contain more precise boundaries and finer details.
+Specifically, it achieved better performance at distinguishing between visually similar object classes like the road and sidewalk.
+Additional Comparison with Cross-modal Contrastive Methods.
+In Table I, we evaluate the semantic segmentation based on the settings used by ALSO [15].
+However, those settings differ from the ones adopted in contrastive methods such as [16, 17, 18].
+To align with the settings adopted in these contrastive methods, we have conducted additional experiments based on their experimental protocols.
+As presented in Table II, our approach also outperforms these baselines across the nuScenes and SemanticKITTI datasets.
+Specifically, our method shows significant improvements in both linear probing and fine-tuning settings.
+For linear probing with 100% of annotations on nuScenes, our approach achieves 42.81% mIoU, surpassing the previous state-of-the-art ST-SLidR (40.48%).
+When fine-tuning with only 1% of annotations, our method achieves a mIoU of 42.29%, outperforming ST-SLidR (40.75%).
+Furthermore, our method demonstrates strong performance on the SemanticKITTI dataset, reinforcing its generalization capability across different datasets and tasks.
+TABLE III: Comparisons between our method and other self-supervised learning methods fine-tuned on the KITTI3D dataset.
 We report the AP evaluated with 40 recall positions on the val set of the KITTI3D dataset.
 Note ProposalContrast [7] is specifically tailored for 3D object detection and pre-trained with the large-scale Waymo dataset.
+The best results are highlighted in bold.
 Detector Pre-training Schedule Car Pedestrian Cyclist Avg
@@
 NCLR (Ours) 92.43 84.86 82.58 67.89 60.48 55.47 92.45 74.05 70.29 75.61
-Qualitative Results.
-Fig.
-5 illustrates qualitative results derived from various methods.
-Our self-supervised pre-training method produced more accurate segmentation results compared to other self-supervised learning approaches.
-The segmentation results from our method contain more precise boundaries and finer details.
-Specifically, it achieved better performance at distinguishing between visually similar object classes like the road and sidewalk.
-TABLE III: The results (mAP) for models fine-tuned with varying quantities of annotated data.
+IV-C 3D Object Detection
+Network Architectures.
+In the downstream 3D object detection task, we explore two prevalent network architectures: the SECOND [44] and PV-RCNN [45] object detectors.
+Both architectures are built upon a shared foundational design, which includes a 3D sparse encoder.
+This encoder processes the input voxels through 3D sparse convolutional operations.
+In addition, they incorporate a bird’s-eye-view encoder (termed the 2D-backbone), which is activated post-BEV projection.
+The key difference between them lies in their respective detection heads.
+SECOND employs a region proposal network (RPN) directly on the 2D backbone, whereas PV-RCNN enhances the RPN predictions through point-level refinement.
+This refinement leads to more accurately defined bounding boxes and improved confidence estimations.
+Pre-training Protocol.
+In alignment with the requirements of downstream detection tasks, we process the point clouds through a voxelization step.
+This involves setting the grid size to [0.05m, 0.05m, 0.1m] along the $x,y$ and $z$ axes respectively.
+The maximum number of non-empty input voxels is limited to 60,000.
+The raw image is resized to (512, 160) as input.
+We utilize the default AdamW optimizer.
+The settings of the optimizer include a peak learning rate of 0.001 and a weight decay factor of 0.001.
+Cosine learning rate schedule [53] is adopted.
+We pre-train the backbone for 50 epochs on the SemanticKITTI dataset with a batch size of 8 on a single GPU.
+Regarding the configuration of the VoxelNet backbone [43], we ensure consistency by employing parameters identical to those used in the corresponding 3D object detectors.
+TABLE IV: The results (mAP) for models fine-tuned with varying quantities of annotated data.
 These results specifically reflect the performance in 3D object detection under moderate difficulty conditions on the validation set of the KITTI3D dataset.
@@
 NCLR (Ours) 82.95 58.91 73.67 71.84(+1.95)
-4.3 3D Object Detection
-Network Architectures.
-In the downstream 3D object detection task, we explore two prevalent network architectures: the SECOND [44] and PV-RCNN [45] object detectors.
-Both architectures are built upon a shared foundational design, which includes a 3D sparse encoder.
-This encoder processes the input voxels through 3D sparse convolutional operations.
-In addition, they incorporate a bird’s-eye-view encoder (termed the 2D-backbone), which is activated post BEV projection.
-The key difference between them lies in their respective detection heads.
-SECOND employs a region proposal network (RPN) directly on the 2D backbone, whereas PV-RCNN enhances the RPN predictions through point-level refinement.
-This refinement leads to more accurately defined bounding boxes and improved confidence estimations.
-Pre-training Protocol.
-In alignment with the requirements of downstream detection tasks, we process the point clouds through a voxelization step.
-This involves setting the grid size to [0.05m, 0.05m, 0.1m] along the $x,y$ and $z$ axes respectively.
-The maximum number of non-empty input voxels is limited to 60,000.
-The raw image is resized to (512, 160) as input.
-We utilize the default AdamW optimizer.
-The settings of the optimizer include a peak learning rate of 0.001 and a weight decay factor of 0.001.
-Cosine learning rate schedule [53] is adopted.
-We pretrain the backbone for 50 epochs on SemanticKITTI dataset with a batch size of 8 on a single GPU.
-Regarding the configuration of the VoxelNet backbone [43], we ensure consistency by employing parameters identical to those used in the corresponding 3D object detectors.
+TABLE V: Comparisons between our method and other pre-training methods by finetuning on 5% of the Waymo 3D detection dataset.
+We report the mAP and mAPH metrics at LEVEL_2 on the validation set.
+Random initialization denotes training from scratch.
+The best results are highlighted in bold.
+Initialization Overall Vehicle Pedestrian Cyclist
+mAP mAPH mAP mAPH mAP mAPH mAP mAPH
+Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
+PointContrast [3] 45.32 41.30 52.12 51.61 53.68 43.22 30.16 29.09
+ProposalContrast [7] 46.62 42.58 52.67 52.19 54.31 43.82 32.87 31.72
+MV-JAR [54] 50.52 46.68 56.47 56.01 57.65 47.69 37.44 36.33
+PonderV2 [37] 50.87 46.81 56.54 55.49 58.09 48.41 37.98 36.52
+Ours 51.86 48.32 57.13 55.93 59.32 50.15 39.14 38.89
 Downstream Fine-tuning Protocol.
@@
 Consistent with the methodology outlined in a prior study [15], this fine-tuning process is repeated three times independently, and the highest mean Average Precision (mAP) achieved on KITTI3D’s validation set is recorded and presented.
-TABLE IV: The performance metrics (PQ and IoU) for panoptic segmentation on SemanticKITTI showcase the effectiveness of the pre-trained models when fine-tuning with various proportions of annotated data.
+TABLE VI: The performance metrics (PQ and IoU) for panoptic segmentation on SemanticKITTI showcase the effectiveness of the pre-trained models when fine-tuning with various proportions of annotated data.
+The best results are highlighted in bold.
 Pre-training Schedule 0.10%   1%   10% 50% 100%
@@
 NCLR (Ours) 12.79 19.14 30.69 36.24 52.79 58.78 56.78 62.54 56.87 62.69
-TABLE V: Experimental results from models pre-trained through self-supervised and supervised training on SemanticKITTI, subsequently fine-tuned for panoptic segmentation tasks using both the mini and complete training sets of nuScenes.
+Quantitative Evaluation Results.
+We evaluate the transferability of our pre-trained model by first pre-training on the SemanticKITTI dataset and subsequently fine-tuning on the KITTI3D dataset.
+We report the experimental results in Table III.
+It is evident that our approach consistently surpasses the baseline established by training from scratch, achieving an enhancement for both SECOND and PV-RCNN.
+With our pre-training method, the performance improves, particularly in the Pedestrian and Cyclist categories, reaching an average mAP of 70.38%.
+Our method performs comparably to state-of-the-art models like ProposalContrast [7], which is specifically tailored for detection tasks and pre-trained with the large-scale Waymo dataset.
+One of the primary benefits of self-supervised learning lies in its ability to improve the performance of downstream tasks when the annotation source is limited.
+To further assess this aspect, we evaluated our method in label-efficient 3D object detection.
+We consider a model trained from a state of random initialization as our standard for comparison.
+The superiority of our pre-trained model compared to this standard is detailed in Table IV.
+Essentially, our pre-trained model consistently boosts the performance in detection tasks using both SECOND and PV-RCNN architectures, particularly notable under conditions of limited labeled data – showing an improvement of 2.43% and 2.89% in mAP with just 1% labeled data for SECOND and PV-RCNN, respectively.
+Besides, our model surpasses the baseline performances across all variations of available annotated data quantities.
+We compare NCLR with MAE-based MV-JAR [54] and rendering-based PonderV2 [37] on the Waymo 3D detection benchmark.
+Specifically, we employ the official implementation of MV-JAR [54] and SST [55] for pre-training and downstream evaluation, respectively.
+As shown in Table V, our method consistently outperforms the MAE-based MV-JAR [54] and PonderV2 [37] across most evaluation metrics on the Waymo 3D detection dataset when fine-tuning with 5% of the available annotations.
+Specifically, our method achieves the best overall results, i.e., 51.86% mAP and 48.32% mAPH.
+It represents improvements of +0.99% and +1.51%, respectively, compared to the rendering-based baseline, PonderV2 [37].
+TABLE VII: Experimental results from models pre-trained through self-supervised and supervised training on SemanticKITTI, subsequently fine-tuned for panoptic segmentation tasks using both the mini and complete training sets of nuScenes.
 We report the PQ and IoU metrics on the full validation set of nuScenes.
@@
 NCLR (Ours) 33.37 31.62 53.97 59.76
-Quantitative Evaluation Results.
-We evaluate the transferability of our pre-trained model by first pre-training on the SemanticKITTI dataset and subsequently fine-tuning on the KITTI3D dataset.
-We report the experimental results in Table II.
-It is evident that our approach consistently surpasses the baseline established by training from scratch, achieving an enhancement for both SECOND and PV-RCNN.
-With our pre-training method, the performance improves, particularly in the Pedestrian and Cyclist categories, reaching an average mAP of 70.38%.
-Our method performs comparably to state-of-the-art models like ProposalContrast [7], which is specifically tailored for detection tasks and pre-trained with the large-scale Waymo dataset.
-One of the primary benefits of self-supervised learning lies in its ability to improve the performance of downstream tasks when the annotation source is limited.
-To further assess this aspect, we evaluated our method in label-efficient 3D object detection.
-We consider a model trained from a state of random initialization as our standard for comparison.
-The superiority of our pre-trained model compared to this standard is detailed in Table III.
-Essentially, our pre-trained model consistently boosts the performance in detection tasks using both SECOND and PV-RCNN architectures, particularly notable under conditions of limited labeled data – showing an improvement of 2.43% and 2.89% in mAP with just 1% labeled data for SECOND and PV-RCNN, respectively.
-Besides, our model surpasses the baseline performances across all variations of available annotated data quantities.
-4.4 Panoptic Segmentation
+IV-D Panoptic Segmentation
 To assess the instance-level features learned by our method, we further fine-tune our pre-trained models for the task of panoptic segmentation.
@@
 For the clustering phase, we opt for the mean shift algorithm  [56] with a set bandwidth of 1.2 and a minimum cluster size requirement of 50 points.
-The other architectures and settings are the same as Section 4.2 during pre-training.
+The other architectures and settings are the same as Section IV-B during pre-training.
 Protocol for Fine-tuning on Downstream Tasks.
@@
 We pre-train the backbone on the SemanticKITTI dataset and fine-tune it on different percentage subsets of SemanticKITTI dataset.
-As shown in Table IV, our method is consistently better than previous self-supervised pre-training approaches.
+As shown in Table VI, our method is consistently better than previous self-supervised pre-training approaches.
 When the segmentation model is trained with fewer labels, we can observe an obvious improvement compared to the baseline of scratch training.
@@
 We initially pre-train the 3D backbone on the SemantiKITTI dataset and subsequently fine-tune it on both the full and mini training sets of nuScenes.
-As presented in Table V, while all approaches enhance performance on the nuScenes dataset, the features generated through our technique demonstrate greater adaptability when employed in a different domain dataset.
+As presented in Table VII, while all approaches enhance performance on the nuScenes dataset, the features generated through our technique demonstrate greater adaptability when employed in a different domain dataset.
 Furthermore, in comparison to supervised pre-training on SemanticKITTI, our approach exhibits improved results.
@@
 These similarity maps are illustrated using examples from the SemanticKITTI validation set.
-4.5 Ablation Studies
+IV-E Ablation Studies
 In our ablation studies on the nuScenes dataset, we adhere to the evaluation protocol from ALSO [15].
@@
 We limit the training period for these studies to 100 epochs.
-TABLE VI: Ablation study on nuScenes semantic segmentation. End-to-end pose estimation indicates using soft-matching, differentiable EPnP solver, and loss of pose estimation.
+TABLE VIII: Ablation study on nuScenes semantic segmentation. End-to-end pose estimation indicates using soft-matching, differentiable EPnP solver, and loss of pose estimation.
 Baseline
@@
 Effect of Key Components.
-We analyze the impact of each component incorporated into our method in Table VI,.
+We analyze the impact of each component incorporated into our method in Table VIII,.
 We can observe that our proposed key components benefit the pre-training framework in learning useful representation and yield better performance in downstream tasks.
@@
 Then we analyze the effect of learnable transformation alignment in detail.
-TABLE VII: Comparison of similarities of image and corresponding point features learned w/o and w/ learnable transformation alignment (LTA).
+TABLE IX: Comparison of similarities of image and corresponding point features learned w/o and w/ learnable transformation alignment (LTA).
 We measure the mean and standard value of similarities on the validation set of SemanticKITTI.
@@
 7.
-We also provide the quantitative comparison of similarities between features of two modalities in Table VII.
+We also provide the quantitative comparison of similarities between features of two modalities in Table IX.
 Specifically, we compute the cosine similarity between the features of the points and corresponding pixels with and without linear feature transformation, respectively.
@@
 However, applying the learnable feature transformation bridges the gap between the two modalities and brings more reasonable similarities.
-TABLE VIII: Feature matching accuracy on SemanticKITTI dataset.
+TABLE X: Feature matching accuracy on SemanticKITTI dataset.
 The error indicates the discrepancy between the locations of matched pixels and the actual projections of points on the image plane.
@@
 Learnable + Soft-Match 2.34 $\pm$ 1.05 90.28
-4.6 Feature Matching Accuracy
+IV-F Feature Matching Accuracy
 We present a comparative analysis of feature matching accuracy using various feature alignment and point-to-pixel matching methods.
-The mean and standard deviation of error distances are detailed in Table VIII.
+The mean and standard deviation of error distances are detailed in Table X.
 To assess the precision of different methods, we calculate the matching accuracy (Acc.), defined as the percentage of fine matchings with errors less than 5 pixels.
-Results in Table VIII demonstrate that the integration of learnable transformation alignment and the soft-matching approach enhances matching accuracy on the SemanticKITTI dataset.
+Results in Table X demonstrate that the integration of learnable transformation alignment and the soft-matching approach enhances matching accuracy on the SemanticKITTI dataset.
 Consequently, our method facilitates the network’s comprehensive understanding of the relationship between image and point cloud data, learning effective representation to process information from each domain.
@@
 The $x$-axis represents the threshold levels for RTE and RRE, while the $y$-axis indicates the recall rate, reflecting the proportion of samples where RREs or RTEs remain below the corresponding threshold.
+TABLE XI: Comparison of registration accuracy (mean ± std) achieved by various methods on the SemanticKITTI and nuScenes datasets.
+The metric “RTE” measured in meters and “RRE” measured in degrees are used for evaluation.
+The best results are highlighted in bold.
+Method SemanticKITTI nuScenes
+RTE $\downarrow$ (m) RRE $\downarrow$ () RTE $\downarrow$ (m) RRE $\downarrow$ ()
+BANet [57]+ICP $3.47\pm 1.44$ $5.21\pm 3.76$ $4.09\pm 2.28$ $7.85\pm 3.62$
+Grid Cls. + EPnP  [58] $1.07\pm 0.61$ $6.48\pm 1.66$ $2.35\pm 1.12$ $7.20\pm 1.65$
+DeepI2P (3D) [58] $1.27\pm 0.80$ $6.26\pm 2.29$ $2.00\pm 1.08$ $7.18\pm 1.92$
+DeepI2P (2D) [58] $1.46\pm 0.96$ $4.27\pm 2.74$ $2.19\pm 1.16$ $3.54\pm 2.51$
+EFGHNe [59] $3.19\pm 1.13$ $4.95\pm 2.50$ $3.92\pm 1.49$ $5.74\pm 3.44$
+CorrI2P [49] $0.74\pm 0.65$ $2.07\pm 1.64$ $1.83\pm 1.06$ $2.65\pm 1.93$
+VP2P [60] $0.67\pm 0.48$ $2.43\pm 4.47$ $0.89\pm 1.44$ $2.15\pm 7.03$
+Ours $\boldsymbol{0.27\pm 0.26}$ $\boldsymbol{1.45\pm 1.04}$ $\boldsymbol{0.72\pm 0.63}$ $\boldsymbol{1.25\pm 0.93}$
 Figure 10: Visual results of detecting overlapping areas between images and point clouds.
@@
 In the first columns, we show the serious misalignment between the image and point cloud, which has initially undergone random rotation and translation.
-In the second column, we project the point cloud on the image plane with our estimated spatial transformation between the camera and LiDAR systems.
-Meanwhile, the third column illustrates the projections of the point cloud with ground-truth rigid transformation (camera pose).
+In the second column, we project the point cloud on the image plane with our estimated spatial pose aligning the camera and LiDAR systems.
+Meanwhile, the third column illustrates the projections of the point cloud with ground-truth rigid pose (camera pose).
 The painted color indicates the depth of each LiDAR point.
-4.7 Accuracy of Image-to-Point Cloud Calibration
+IV-G Accuracy of Image-to-Point Cloud Calibration
 We compared our approach with other leading methods for image-to-point cloud registration/calibration [58, 49, 60].
@@
 The results clearly indicate the enhanced performance of our technique in terms of registration accuracy.
+We also report the quantitative results presented in Table XI to highlight the effectiveness of our proposed method in 2D-3D calibration across the SemanticKITTI and nuScenes datasets.
+Specifically, our method outperforms all baseline approaches by achieving the lowest RTE and RRE on both datasets.
+On the SemanticKITTI dataset, our method achieves an RTE of 0.27 ± 0.26 meters and an RRE of 1.45 ± 1.04 degrees, representing substantial improvements over the competing methods.
+Notably, the closest baseline, VP2P, reports an RTE of 0.67 ± 0.48 meters and an RRE of 2.43 ± 4.47 degrees, but our approach cuts down the translation error by over 50% while reducing the rotation error by nearly 40%.
+On the nuScenes dataset, our method maintains its superior performance with an RTE of 0.72 ± 0.63 meters and an RRE of 1.25 ± 0.93 degrees.
+The consistent reduction in both RTE and RRE demonstrates our model’s robust ability to accurately estimate the rigid pose between the camera and LiDAR systems across different datasets and scenarios.
 While numerous studies have addressed the task of image-to-point cloud registration, it is noteworthy that these methods predominantly utilize simulated data.
@@
 In our research, we broaden the scope of image-to-point cloud registration/calibration by extending its application to self-supervised pre-training for 3D perception tasks.
-4.8 Further Discussions
+IV-H Further Discussions
 Fig.
@@
 The figures in first column reveal noticeable misalignments between the image and the point cloud initialized with random rotation and translation.
-Despite the challenging initial conditions, our method accurately estimates the applied rigid transformation.
-This precise detection of overlapping areas and accurate estimation of rigid transformations indicate that the network could effectively understand the information and relationships between each domain.
+Despite the challenging initial conditions, our method accurately estimates the applied rigid pose.
+This precise detection of overlapping areas and accurate estimation of the rigid pose indicate that the network effectively understands the spatial information and relationships between each domain.
 Such insights help explain why the representations learned in the pretext task are advantageous for downstream tasks.
-5 Conclusion
+V Conclusion
 We have introduced an innovative perspective to self-supervised learning by achieving a thorough alignment between two distinct modalities.